OpenStudy (anonymous):

2x-5y=10 (intercepts)

OpenStudy (anonymous):

for x - intercept put y = 0 and for y - intercept put x = 0

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

2x-5y=10 Or 2x/10 -5y/10=1 Or x/5+y/-2 =1 X-intercept=5 Y-intercept=-2

OpenStudy (anonymous):

Or how about just: with x=0, then: -5y=10 -y=2 y=-2 with y=0 then: 2x=10 x=5